# STEMS: Spatial-Temporal Enhanced Safe Multi-Agent Coordination for Building Energy Management

### v1

@@ IV-A Building Energy Management Systems
Building energy management systems play a critical role in achieving global carbon neutrality goals, as buildings account for approximately 40% of worldwide energy consumption. Traditional approaches have primarily focused on individual building optimization through three main categories: rule-based control systems that use predefined logic for equipment operation, MPC methods that optimize future actions based on system models [2, 18], and heuristic algorithms that provide practical solutions for complex optimization problems[19].
The integration of renewable energy sources introduces substantial uncertainty due to their intermittent nature, while dynamic electricity pricing and unpredictable occupancy patterns create complex optimization landscapes [1]. These uncertainties make it difficult for traditional model-based approaches to maintain optimal performance. Furthermore, conventional methods typically adopt a single-building optimization perspective, failing to account for the interdependencies and coordination opportunities that exist among multiple buildings in communities or districts. Recent advances in deep learning have led to the emergence of RL based approaches for building energy management [10, 17, 12]. RL methods offer model-free learning capabilities that can adapt to system uncertainties without requiring detailed domain knowledge. Most existing RL applications focus on individual building control due to the inherent complexity and stability concerns of multi-agent systems. This single-building focus leads to suboptimal performance in real-world implementations, where multiple buildings within a community pursue their local objectives independently, often resulting in competitive behaviors that prevent global optimization[9].
@@ IV-B Multi-Agent Reinforcement Learning
Multi-agent reinforcement learning provides a principled framework for coordinated decision-making in distributed systems where multiple autonomous agents interact within a shared environment [16, 20, 21]. The core challenge lies in handling non-stationarity, where each agent’s environment changes as other agents adapt their policies, creating a moving target problem. Classical multi-agent RL algorithms address this through centralized training with decentralized execution methods like MADDPG [16], value function factorization techniques such as QMIX [20], and policy gradient methods that account for multi-agent interactions. Building energy management represents a natural application domain for multi-agent RL due to the distributed nature of building systems and their inherent interdependencies. Distributed residential energy management systems employ multi-agent RL to coordinate household energy scheduling, where each building optimizes its energy consumption while contributing to grid stability [3, 22]. Multi-zone HVAC control leverages multi-agent RL for temperature regulation across building zones, handling occupancy patterns while maintaining thermal comfort [13]. Grid-interactive building coordination uses iterative sequential action selection for effective load shaping and demand response [12]. Federated learning approaches have also emerged to enable collaborative optimization while preserving privacy [23]. However, current multi-agent RL methods lack the ability to effectively model spatial-temporal dependencies among buildings, which limits coordination effectiveness. Spatial dependencies arise from physical coupling between buildings, including thermal transfer, shading effects, and shared infrastructure. Temporal dependencies encompass both short-term dynamics, such as equipment response times and occupancy patterns, and long-term trends including seasonal variations. Graph neural networks have shown promise in capturing spatial relationships in power grids and building clusters [5, 4]. Effective modeling of these spatial-temporal relationships through graph-based representations could enable more informed coordination decisions and enhanced information propagation among agents.
@@ IV-C Safe Reinforcement Learning
Safety is a critical concern when deploying RL algorithms in real-world systems, particularly in safety-critical domains such as building energy management [8, 24]. Current safe RL approaches can be categorized into several main directions: constrained optimization methods that integrate safety requirements directly into the learning objective, CBFs that provide mathematical guarantees for constraint satisfaction in continuous control systems, and safe exploration strategies that prevent dangerous actions during the learning process [25]. Recent work has demonstrated the application of safe RL to smart home energy management, where safety constraints include battery state-of-charge limits, temperature comfort bounds, and power consumption limits [7, 6, 11]. However, most existing safe RL methods focus on single-agent scenarios and lack comprehensive frameworks for multi-agent coordination with safety guarantees. The integration of safety constraints with multi-agent coordination presents unique challenges in building energy systems. Distributed safety constraints must be satisfied not only at the individual building level but also at the community and grid levels. Recent surveys highlight that safe RL for power systems remains an active research area with significant gaps in multi-agent scenarios [26]. The combination of spatial-temporal awareness with rigorous safety guarantees in multi-agent building energy management represents a key research direction that requires novel approaches to ensure both coordination effectiveness and safety compliance.
@@ V Conclusion
In this work, we study the coordinated building energy management problem in multi-building systems with complex spatial-temporal dependencies and safety requirements. We propose STEMS, a safety-constrained multi-agent reinforcement learning framework for coordinated building energy management based on information sharing and cooperative optimization. The GCN-Transformer fusion architecture is introduced to multi-agent RL to enhance agents’ spatial-temporal awareness of inter-building relationships and temporal energy patterns. The CBF-based safety shield provides mathematical guarantees for safe operation while maintaining system performance. In experiments, we verify the effectiveness and advantages of our method and each of its components in both safety and efficiency by comparing results with baseline methods across different scenarios, including long-term performance evaluation and extreme weather conditions with diverse building types. The results demonstrate significant improvements in cost, emission and safety violation reduction while maintaining optimal comfort levels. Future work could extend to enhance the robustness of the multi-agent RL algorithm and CBF safety mechanisms with different types of uncertain renewable energy generation, evaluate scalability on larger-scale building networks, and explore the interpretability of learned policies to provide practical insights for building management applications.

### v2

@@ IV-A Building Energy Management Systems
Building energy management systems play a critical role in achieving global carbon neutrality goals, as buildings account for approximately 40% of worldwide energy consumption. Traditional approaches have primarily focused on individual building optimization through three main categories: rule-based control systems that use predefined logic for equipment operation, MPC methods that optimize future actions based on system models [zhang2022building-survey, berberich2024overview], and heuristic algorithms that provide practical solutions for complex optimization problems[ghalambaz2021building].
The integration of renewable energy sources introduces substantial uncertainty due to their intermittent nature, while dynamic electricity pricing and unpredictable occupancy patterns create complex optimization landscapes [chen2024green]. These uncertainties make it difficult for traditional model-based approaches to maintain optimal performance. Furthermore, conventional methods typically adopt a single-building optimization perspective, failing to account for the interdependencies and coordination opportunities that exist among multiple buildings in communities or districts. Recent advances in deep learning have led to the emergence of RL based approaches for building energy management [guo2022real, MetaEMS, MARLISA]. RL methods offer model-free learning capabilities that can adapt to system uncertainties without requiring detailed domain knowledge. Most existing RL applications focus on individual building control due to the inherent complexity and stability concerns of multi-agent systems. Recent single-agent approaches such as CLUE[safeMBRL] employ GP-based methods with meta-kernel learning for HVAC control, achieving data efficiency through transfer learning from reference buildings. However, the GP-based approach involves $O(n^{3})$ computational complexity that scales with dataset size, limiting scalability for long-term control in dynamic multi-building environments. This single-building focus leads to suboptimal performance in real-world implementations, where multiple buildings within a community pursue their local objectives independently, often resulting in competitive behaviors that prevent global optimization[meng2024online].
@@ IV-B Multi-Agent Reinforcement Learning
Multi-agent reinforcement learning provides a principled framework for coordinated decision-making in distributed systems where multiple autonomous agents interact within a shared environment [lowe2017multi-maddpg, Rashid2018QMIX, yu2022surprising]. The core challenge lies in handling non-stationarity, where each agent’s environment changes as other agents adapt their policies, creating a moving target problem. Classical multi-agent RL algorithms address this through centralized training with decentralized execution methods like MADDPG [lowe2017multi-maddpg], value function factorization techniques such as QMIX [Rashid2018QMIX], and policy gradient methods that account for multi-agent interactions. Building energy management represents a natural application domain for multi-agent RL due to the distributed nature of building systems and their inherent interdependencies. Distributed residential energy management systems employ multi-agent RL to coordinate household energy scheduling, where each building optimizes its energy consumption while contributing to grid stability [kumari2024multi, ding2024distributed]. Multi-zone HVAC control leverages multi-agent RL for temperature regulation across building zones, handling occupancy patterns while maintaining thermal comfort [MA-HVAC]. Grid-interactive building coordination uses iterative sequential action selection for effective load shaping and demand response [MARLISA]. Federated learning approaches have also emerged to enable collaborative optimization while preserving privacy [xia2025federated]. Recent work such as MAIL[gao2025multi] demonstrates notable training efficiency through imitation learning from MILP solvers, enabling rapid policy learning by leveraging expert demonstrations. However, this approach requires accurate MILP formulation of the problem, which poses challenges for capturing nonlinear thermal dynamics and multi-objective optimization in real-world building systems. Despite these advances, current multi-agent RL methods [saberi2024madcq, tariq2025dmappo] lack the ability to effectively model spatial-temporal dependencies among buildings, which limits coordination effectiveness. Spatial dependencies arise from physical coupling between buildings, including thermal transfer, shading effects, and shared infrastructure. Temporal dependencies encompass both short-term dynamics, such as equipment response times and occupancy patterns, and long-term trends including seasonal variations. Graph neural networks have shown promise in capturing spatial relationships in power grids and building clusters [chen2023graph, haidar2023selective, dmg, knnmts, LMHR]. Effective modeling of these spatial-temporal relationships through graph-based representations could enable more informed coordination decisions and enhanced information propagation among agents.
@@ IV-C Safe Reinforcement Learning
Safety is a critical concern when deploying RL algorithms in real-world systems, particularly in safety-critical domains such as building energy management [brunke2022safe, gu2024review]. Current safe RL approaches can be categorized into several main directions: constrained optimization methods that integrate safety requirements directly into the learning objective, CBFs that provide mathematical guarantees for constraint satisfaction in continuous control systems, and safe exploration strategies that prevent dangerous actions during the learning process [thomas2021safe]. Recent work has demonstrated the application of safe RL to smart home energy management, where safety constraints include battery state-of-charge limits, temperature comfort bounds, and power consumption limits [ding2022safe, wang2025safe, Jang_Yan_Spangher_Spanos_2024]. However, most existing safe RL methods focus on single-agent scenarios and lack comprehensive frameworks for multi-agent coordination with safety guarantees. The integration of safety constraints with multi-agent coordination presents unique challenges in building energy systems. Distributed safety constraints must be satisfied not only at the individual building level but also at the community and grid levels. Recent surveys highlight that safe RL for power systems remains an active research area with significant gaps in multi-agent scenarios [su2025review]. The combination of spatial-temporal awareness with rigorous safety guarantees in multi-agent building energy management represents a key research direction that requires novel approaches to ensure both coordination effectiveness and safety compliance.
@@ V Conclusion
In this work, we study the coordinated building energy management problem in multi-building systems with complex spatial-temporal dependencies and safety requirements. We propose STEMS, a safety-constrained multi-agent reinforcement learning framework for coordinated building energy management based on information sharing and cooperative optimization. The GCN-Transformer fusion architecture is introduced to multi-agent RL to enhance agents’ spatial-temporal awareness of inter-building relationships and temporal energy patterns. The CBF-based safety shield provides mathematical guarantees for safe operation while maintaining system performance. In experiments, we verify the effectiveness and advantages of our method and each of its components in both safety and efficiency by comparing results with baseline methods across different scenarios, including long-term performance evaluation and extreme weather conditions with diverse building types. The results demonstrate significant improvements in cost, emission and safety violation reduction while maintaining optimal comfort levels. Future work could extend to enhance the robustness of the multi-agent RL algorithm and CBF safety mechanisms with different types of uncertain renewable energy generation, incorporate outage-aware training to further enhance resilience, and scale STEMS to large-scale building networks while maintaining real-time performance guarantees.